OpenStudy (anonymous):

Anyone in Algebra 2 that goes to CCA? I'm 21 lesson days behind and I NEED help, I would love to be able to graduate my senior year (this year)...

OpenStudy (whpalmer4):

Start asking questions, and we'll start helping you.

OpenStudy (anonymous):

Warning: I don't understand how to do ANY of it... Here's a few problems I need help with. 1. describe the horizontal asymptote(s) and hole(s) for the graph of y=x-4/x^2+3x+2 2. find the horizontal asymptote of the graph of y=-2x^6+5x+8/ 8x^6+6x+5 3.Simplify the rational expression. State any restrictions on the variable. t^2+2t-24/t^2-36

OpenStudy (whpalmer4):

Do you understand what a horizontal asymptote is? How about a hole?

OpenStudy (anonymous):

nope. I honestly don't

OpenStudy (whpalmer4):

Okay, let's see if we can fix that. The easiest thing to understand (at least I think so) is the concept of a vertical asymptote and restrictions on a variable. Looking at the 3rd problem: \[\frac{t^2+2t-24}{t^2-36}\]Can you factor the numerator and denominator for me?

OpenStudy (anonymous):

sure if I knew what that meant.

OpenStudy (whpalmer4):

Do you understand how to factor polynomials? It's like undoing multiplication.

OpenStudy (whpalmer4):

For example, going from \[x^2+2x+1\]to\[(x+1)(x+1)\]because \[(x+1)(x+1) = x*x + 1*x + 1*x + 1*1 = x^2 + 2x + 1\]

OpenStudy (whpalmer4):

Factoring is a necessary skill to do these problems, so if you don't know how to do it, we'll have to back up and you'll have to learn it.

OpenStudy (anonymous):

The thing is I took the first semester of Algebra 2 my Junior year and I was sucking badly so then I asked my counselor to drop it and just put me in a full class of trigonometry my senior year (this year) and she said she would but she didn't so she left it on my transcript as a D and I honestly have absolutely NO idea what's going on and it's not because I'm stupid because I was really good at pre algebra, algebra 1, and geometry when I was in regular school, but online its just harder and I don't understand ANYTHING. It's seriously depressing me because I am very afraid I won't graduate....

OpenStudy (whpalmer4):

I'm certainly sympathetic to the difficulties one might encounter with online education. One of them is that you might not have a teacher with whom you can interact. However, today you have one, and he's trying to interact with you by asking you questions to gain a better understanding of where the gaps in your knowledge may be, and at least one of them has gone unanswered. So, again, do you, or do you not know how to factor polynomials? I'm not trying to make you feel bad about your lack of knowledge; I need to know where we need to start. You can't build a 70 story skyscraper if you skip building the 3rd floor, right?

OpenStudy (anonymous):

Okay well no I have forgotten how to factor polynomials

OpenStudy (whpalmer4):

But you think you knew at one point? It should come back pretty quickly, if so. Might be a slower process if you're learning it for the first time.

OpenStudy (whpalmer4):

Let's take the example I used earlier: \[x^2+2x+1\]We want to factor that into the product of two binomials: \[x^2+2x+1 = (x+a)(x+b)\]What needs to be true for that to happen? Let's multiply \((x+a)(x+b)\) and see what it looks like. \[(x+a)(x+b) = x(x+b) + a(x+b)\]by the distributive property, which we'll apply again: \[(x+a)(x+b) = x*x + x*b + a*x + a*b\]now we collect like terms \[(x+a)(x+b) = x^2 + (a+b)x + a*b\] Any question about that so far?

OpenStudy (anonymous):

Yes. I do not understand how you got (x+a)(x+b)= x(x+b) + a(x+b)

OpenStudy (whpalmer4):

Distributive property: a(b+c) = a*b + a*c

OpenStudy (anonymous):

You know if you're just going to send me rude messages then you can quit "helping me" I'm sorry if I like to take my time to try to figure out the things you're teaching me, this isn't a fast learning process for everyone.

OpenStudy (whpalmer4):

I can see when you are looking at other questions. You are the one being rude by not paying attention to the person trying to help you.

OpenStudy (anonymous):

Actually for your information I was paying attention to you I had two different screens up at once.